# The Atacama Cosmology Telescope: Quantifying Atmospheric Emission above Cerro Toco

### v1

@@ Abstract
At frequencies below 1 Hz, fluctuations in atmospheric emission in the Chajnantor region in northern Chile are the primary source of interference for bolometric millimeter-wave observations. This paper focuses on the statistics of these fluctuations using measurements from the Atacama Cosmology Telescope (ACT) and the Atacama Pathfinder Experiment (APEX) water vapor radiometer. After introducing a method for separating atmospheric effects from other systematic effects, we present a direct measurement of the temporal outer scale of turbulence of $\tau_{0}\approx 50$s corresponding to a spatial scale of $L_{0}\approx 500$m. At smaller scales, the fluctuations are well described by the Kolmogorov 2/3 power law until, at yet smaller scales, the effects of beam smearing become important. As a part of this study, we present measurements of the atmosphere by the APEX radiometer over 20 years, focused on fluctuations in precipitable water vapor (PWV). We find that the 30-minute mean of the total PWV is not in general a robust estimator of the level of fluctuations. We show that the microwave frequency spectrum of these fluctuations is in good agreement with predictions by the am code for bands above 90 GHz. We then show that the variance of fluctuations in ACT’s mm-wave bands correlates with the variance of fluctuations in PWV measured by APEX, even though the observatories are 6 km apart and observe different lines of sight. We find that ACT’s atmosphere-determined optical efficiencies are consistent with previous planet-based results.
@@ 2 Atmospheric conditions in the Chajnantor region
Water vapor fluctuates on short timescales due to turbulent mixing. Consider a turbulent vortex spanning a gradient in humidity, for example between a wetter layer of air closer to the ground and a dryer layer farther up in the atmosphere. The velocity field will transport air from the wetter later to the dryer layer and vice versa, leading to turbulent variations in water vapor content. This is depicted in Figure 1 and described in Tatarski (1961) as turbulent diffusion.
The degree of turbulence is proportional to the (primarily vertical) gradient of the quantity of interest. The quantities with appreciable gradients are wind velocity, water vapor content compared to the mean fraction and in some cases temperature variation compared to the mean vertical profile, such as when there is a thermal inversion layer [^4] In absence of a temperature inversion, the typical gradient is $-6.5^{\circ}$/km.. Oxygen and nitrogen have comparatively smaller gradients compared to nominal composition. Because of its strong emission lines, water vapor fluctuations are particularly important for mm-wave measurements, but we can think of them as a marker for the turbulence that affects all components of the atmosphere. In the language of Tatarski (1961), water vapor is a “passive additive.” Obukhov (1949) shows that temperature fluctuations can also be treated in the framework of a passive additive.
@@ 3.1 Turbulent distributions as seen by an observer on the ground
To a detector observing the sky, fluctuations in emission will appear as the integral along the line-of-sight through the three-dimensional distribution described above. The total power in each detector is determined by integrating over the contribution from each element in the beam. Under the assumption that the atmosphere is optically thin and the emission is proportional to the water vapor content of the atmosphere, we can, for notational simplicity, represent the total factor of the response of detector $i$ to the atmosphere as a single gain $g_{i}$ such that
@@ 5.1 The ACT instrument
ACT was [^9] ACT was decommissioned in October 2022 to make way for the Simons Observatory (Ade et al., 2019). located in the Chajnantor region of northern Chile at an altitude of 5145 m above mean sea level (GPS altitude, 5188 m). There have been three generations of receivers on the ACT telescope (Fowler et al., 2007; Swetz et al., 2011). In this paper we analyze data from Advanced ACTPol (AdvACT) (Henderson et al., 2016; Ho et al., 2017; Choi et al., 2018; Li et al., 2018, 2021) which uses NIST-built polarization-sensitive detectors operating at 100 mK (Grace et al., 2014; Datta et al., 2014; Ho et al., 2016). The cryostat holds three independent optics tubes each with a polarized array (PA). The optics tubes may be swapped out.
AdvACT comprised the PA4, PA5, PA6, & PA7 dichroic arrays that took the place of ACTPol’s PA1, PA2, and PA3 arrays (Thornton et al., 2016). As shown in Figure 2 of Murphy et al. (2024), the arrays in each optics tube project onto a hexagon on the sky with an edge-to-edge diameter of about $0.75^{\circ}$. The three optics tubes form a triangle with base of about $2^{\circ}$. The entire telescope scans the arrays with an azimuthal speed of $1.5^{\circ}/$sec.
@@ 5.2 Predicted response to fluctuations in PWV
where the passband, $\tau(\nu)$, is the response of the receiver to a Rayleigh-Jeans (RJ) source and normalized to unity, $\eta$ is the overall efficiency of the telescope and receiver, and $T_{b}(\nu)$ is the brightness temperature of the atmosphere as a function of frequency as shown in Figure C1. To model $T_{b}(\nu)$ we use am (version 13.0, see Paine (2019)). The total power in each band, and its derivative with respect to the $w_{z}$, is given in Table C1 as a function of $w_{z}$ for $\eta=1$.
@@
It is more straightforward to work in units of power as opposed to brightness temperature because the first step in ACT’s calibration procedure is to convert the detector output to picowatts (pW). In response to fluctuations $\delta w_{z}$, assuming we receive a single mode of radiation, we measure
@@ 6.1 Time streams and pre-processing
We consider ACT data taken during the 2017, 2018, and 2019 observing seasons (May 2017 to January 2020). We also consider observations at 30 and 40 GHz from the 2020 and 2021 seasons in Section 7.3. We use the same raw data, cuts, and calibration that go into the mapping pipeline (Aiola et al., 2020; Naess et al., 2020). ACT data typically consists of hour-long periods of observations between detector re-biasing, which are then divided into five approximately 11-minute long chunks of time-ordered data (each referred to as a “TOD”). For each TOD, we down sample the data to 100 Hz using a triangular filter. In analyzing the structure function, we remove a linear trend over multiple concatenated TODs to remove an offset and long-term trends in the data. We emphasize that the cosmological analysis corresponds to data at higher frequencies ($>0.5$ Hz) and smaller spatial scales ($\ell\gtrsim 200$ or $\lesssim 1^{\circ}$) than those used here.
@@ 6.2 Mode separation and TOD selection
Figure 6 shows a typical receiver output for 40 minutes. The common mode signal in all three optics tubes and their bands is from the atmosphere and readily apparent. However, there are often optics tube-common modes that do not correspond to emission from the atmosphere or sky. We attribute these to systematic pickup for each optics tube (hereafter “pickup”). When atmospheric fluctuations are weak, these signals can become comparable to or even dominate atmospheric fluctuations, confounding analysis of the atmosphere.
@@
where $g_{w,i}$ and $g_{p,i}$ are the responses of detector $i$ to the fluctuations in water vapor ($\delta w$) and pickup in optics tube $j$ ($\delta p_{j}$), respectively, and $\epsilon_{i}(t)$ is some residual. The pickup modes are common to an optics tube, with each detector in that tube having a different response to it, [^10] Based on a phenomenological examination of the data, we find that a linear response of each detector to the underlying systematic mode in each optics tube is sufficient. while the atmosphere is common to all three optics tubes. [^11] Fluctuations smaller than ACT’s field of view are negligibly small compared to the common mode. The pickup modes are synchronous with the scan but not fixed on the sky nor consistent with a constant wind speed. With one common atmospheric mode and three optics tube pickup modes, we represent the ($n_{\text{detector}}\times n_{\text{time}}$) matrix of all detector fluctuations $\delta\mathbf{P}$ for each TOD, $t$, as
@@
where $\sigma^{2}_{\text{c}}$ and $\sigma^{2}_{\text{nc}}$ are the variances of the common and non-common modes in each detector’s signal. Distributions of $a^{2}$ for each band are shown in Figure B3. We set the threshold for a detector-TOD being “atmosphere-dominated” as $a^{2}\geq 10$, that is, when the variance of the atmospheric fluctuations is at least 10 times that of systematic fluctuations. We restrict the data to detector-TODs with this designation for quantifying the characteristics of turbulence.
@@ 6.3 Structure functions
We model the atmosphere by assuming a constant horizontal wind velocity but this is complicated by the fact that ACT employs a wide back-and-forth constant-elevation scan. We can mitigate the effect of this scan by considering only the correlation between data at the same pointing angle. The back-and-forth scan has the useful property that for every possible delay $\tau$ much less than the total length of the observation, there are always several pairs of co-located points separated by time $\tau$. Considering only these points allows us to compare ACT data directly to Equation 6.
@@ 6.4 Estimating the coefficient of turbulence and outer scale
We model the fluctuations, $y(t)$, as a zero-mean correlated multivariate normal distribution. If the fluctuations are stationary, the covariance depends only on the delay $\tau=|t-t^{\prime}|$ and is given by Equation 6 as
@@ 7.1 Power spectrum, structure functions and outer scale estimates
Figure 7 shows the power spectra of the data in Figure 6. They agree with expectations after accounting for the effects of beam smearing. Evidence for the outer scale shows up at $f<0.01\,$Hz. Figure 8 shows the structure function for each band, averaged [^15] Because the scaling of the structure functions of different TODs can vary by several orders of magnitude, an explicit average will be most sensitive to TODs in which atmospheric fluctuations are the strongest. In order to probe the average shape of the structure function, we take a weighted average to normalize each structure function with respect to its variance. over all TODs that were designated as atmosphere-dominated ($a^{2}\geq 10)$. The key feature is the flattening for $\tau\gtrsim 100$s, beyond which there are no additional contributions to the variance. For each average structure function, we overlay a best fit of the structure function as predicted by the Matérn covariance and accounting for the beam profile and water vapor scale height (see below) as described in more detail in Appendix A.2. Figure 9 shows the distribution of $\tau_{0}$ from the fits for PA6 f150 which is representative of the other arrays. In the frozen transport approximation the spatial length scale differs from the temporal length scale by a factor of the wind speed. Thus, to find the outer scale $L_{0}$ we multiply by the wind speed derived from ERA5 (Hersbach et al., 2020).
@@
As shown in Morris et al. (2022), there is an excellent agreement between the wind speed derived from ACT with that derived from the absolute humidity weighted velocity profile in ERA5.
We also show the distribution of the angular outer scale, computed by scaling $\tau_{0}$ by the angular velocity of the atmosphere estimated directly from ACT following the method in Morris et al. (2022). We see in Figure 9 that $L_{0}$ is comparable to the scale height of the water vapor. While not surprising, simultaneously accounting for both the scale height and outer scale leads to a somewhat modified form of the predicted structure function shown in Figure 8. We discuss this in Appendix A.1.
@@ 7.2 Comparison of fluctuations in ACT to APEX
We find that both $\langle\delta P^{2}\rangle^{\frac{1}{2}}_{t}$ and $\langle\delta w_{z}^{2}\rangle^{\frac{1}{2}}_{t}$, where the subscript denotes average over a TOD, are distributed log-normally. With simulations, we find that the median of each distribution is the best statistic for comparing them. To calibrate ACT’s fluctuation to APEX, we account for the difference in height between ACT and APEX ($\sigma_{ACT}=0.94\,\sigma_{APEX}$, Section 4). This results in a scaled value for $g$ that we call $g_{s}$. To estimate ACT’s efficiency, $\eta$, we divide $g_{s}$ by the entry for $\partial P/\partial w_{z}$ in Table C1. Because the resolution of APEX is 0.01 mm, we compute the correlation for RMS fluctuation levels of 0.02 mm and above. As this cutoff level is increased, the computed $\eta$ decreases. The results are reported in Table 1. Based on the variance over seasons, the dispersion of the correlation in Figure 11, the bias associated with the cutoff level, and the uncertainty in the passbands, we estimate the net uncertainty to be roughly 20%.
@@
There is an overall degeneracy, that we cannot determine, between $\eta$ and any out-of-band passband leakage or an incorrect passband measurement. The quoted values are for the recommended frequency bounds for the integrals over the passbands (Hasselfield et al., 2024) and is based on an analysis of the Fourier Transform Spectrometer (FTS) used to measure them Alford (2024). For the 90 GHz bands, there is possible band pass leakage at the 1% level that is sensitive to the large atmospheric signal near 180 GHz shown in Figure 4. However, there is no clear evidence of sensitivity to this feature.

### v3

@@ Abstract
At frequencies below 1 Hz, fluctuations in atmospheric emission in the Chajnantor region in northern Chile are the primary source of interference for bolometric millimeter-wave observations. This paper focuses on characterizing these fluctuations using measurements from the Atacama Cosmology Telescope (ACT) and the Atacama Pathfinder Experiment (APEX) water vapor radiometer. We show that the total precipitable water vapor (PWV) is not in general an accurate estimator of the level of fluctuations in millimeter-wave atmospheric emission. We also show that the microwave frequency spectrum of atmospheric fluctuations is in good agreement with predictions by the am code for frequency bands above 90 GHz. We introduce a new method for separating atmospheric and systematic fluctuations, allowing us to fit a robust atmospheric flatfield, as well as to study in the atmosphere in greater detail than previous works. We present a direct measurement of the temporal outer scale of turbulence of $\tau_{0}\approx 50$ s corresponding to a spatial scale of $L_{0}\approx 500$ m. Lastly, we show the variance of fluctuations in ACT’s mm-wave bands correlate with the variance of fluctuations in PWV measured by APEX, even though the observatories are 6 km apart and observe different lines of sight.
@@ 2 Atmospheric conditions in the Chajnantor region
Water vapor fluctuates on short timescales due to turbulent mixing. Consider a turbulent vortex spanning a gradient in humidity, for example between a wetter layer of air closer to the ground and a dryer layer farther up in the atmosphere. The velocity field will transport air from the wetter layer to the dryer layer and vice versa, leading to turbulent variations in water vapor content. This is depicted in Figure 1 and described in Tatarski (1961) as turbulent diffusion.
The degree of turbulence is proportional to the (primarily vertical) gradient of the quantity of interest. The quantities with appreciable gradients are wind velocity, water vapor content compared to the mean fraction and in some cases temperature variation compared to the mean vertical profile, such as when there is a thermal inversion layer. [^4] In absence of a temperature inversion, the typical gradient is $-6.5^{\circ}$/km. Oxygen and nitrogen have comparatively smaller gradients compared to nominal composition. Because of its strong emission lines, water vapor fluctuations are particularly important for mm-wave measurements, but we can think of them as a marker for the turbulence that affects all components of the atmosphere. In the language of Tatarski (1961), water vapor is a “passive additive.” Obukhov (1949) shows that temperature fluctuations can also be treated in the framework of a passive additive.
@@ 3.1 Turbulent distributions as seen by an observer on the ground
To a detector observing the sky, fluctuations in emission will appear as the integral along the line-of-sight through the three-dimensional distribution described above. The total power in each detector is determined by integrating over the contribution from each element in the beam. Under the assumption that the atmosphere is optically thin and the emission is proportional to the water vapor content of the atmosphere, we can, for notational simplicity, represent the total factor of the response of detector $i$ to the atmosphere as a single gain $g_{i}$ such that its fluctuation in power $\delta P_{i}$ is given by
@@ 5.1 The ACT instrument
ACT was [^9] ACT was decommissioned in October 2022 to make way for the Simons Observatory (Ade et al., 2019). a 6 m mm-wave telescope located in the Chajnantor region of northern Chile at an altitude of 5145 m above mean sea level (GPS altitude, 5188 m). There were three generations of receivers on the ACT telescope (Fowler et al., 2007; Swetz et al., 2011). In this paper we analyze data from the last, Advanced ACTPol (AdvACT) (Henderson et al., 2016; Ho et al., 2017; Choi et al., 2018; Li et al., 2018, 2021), which used NIST-built polarization-sensitive detectors operating at 100 mK (Grace et al., 2014; Datta et al., 2014; Ho et al., 2016).
For a complete description of the receiver, we refer the reader to (Thornton et al., 2016); here we give the salient features relevant to this work. The roughly 1 m diameter receiver holds three independent, hexagonal, approximately $400$-feed polarization-sensitive detector arrays (PAs), each at the base of a customized “optics tube” that couples the array to the telescope. Looking at the face of the receiver, as shown in (Thornton et al., 2016), these optics tubes are arranged in a triangle with a center typically near $45^{\circ}$ elevation. When projected onto the sky, the combined beam pattern of all three tubes has a triangular pattern with a base of about $2^{\circ}$. At each vertex of the triangle is the beam pattern for one of the hexagonal arrays. The edge-to-edge diameter of the hexagonal pattern is about $0.75^{\circ}$ as shown in Figure 2 of Murphy et al. (2024). The entire telescope scans the arrays with an azimuthal speed of $1.5^{\circ}/$sec.
AdvACT comprised the PA4, PA5, PA6, & PA7 dichroic arrays that took the place of ACTPol’s PA1, PA2, and PA3 (Thornton et al., 2016). The PA4 array measures near 150 and 220 GHz, which we designate as f150 and f220 bands, while the PA5 and PA6 arrays each measure near 90 and 150 GHz, designated as f090 and f150. The PA6 array is highest on the sky. In the last couple of years of ACT’s operation PA6 was swapped out for the lower frequency PA7, operating in the f030 and f040 bands, that is described in Section VII.C below.
@@ 5.2 Predicted response to fluctuations in PWV
where the passband, $\tau(\nu)$, is the response of the receiver to a Rayleigh-Jeans (RJ) source and normalized to unity, $\eta$ is the overall efficiency of the telescope and receiver, and $T_{b}(\nu)$ is the brightness temperature of the atmosphere as a function of frequency as shown in Figure C1. To model $T_{b}(\nu)$ we use am (version 13.0, see Paine (2019)). The total power in each band, and its derivative with respect to the $w_{z}$, is given in Appendix C as a function of $w_{z}$ for $\eta=1$.
@@
It is more straightforward to work in units of power as opposed to brightness temperature because the first step in ACT’s calibration procedure is to convert the detector output to picowatts (pW). In response to fluctuations $\delta w_{z}$, assuming we receive a single mode of radiation, we measure power fluctuation
@@ 6.1 Time streams and pre-processing
We consider ACT data taken during the 2017, 2018, and 2019 observing seasons (May 2017 to January 2020). We also consider observations at 30 and 40 GHz from the 2020 and 2021 seasons in Section 7.3. We use the same raw data, cuts, and calibration that go into the mapping pipeline (Aiola et al., 2020; Naess et al., 2020). ACT data typically consists of hour-long periods of observations between detector re-biasing, which are then divided into five approximately 11-minute long chunks of time-ordered data (each referred to as a “TOD”). For each TOD, we down sample the data from 400 Hz to 100 Hz using a triangular filter (see e.g. (Dünner et al., 2013)). In analyzing the structure function, we remove a linear trend over multiple concatenated TODs to remove an offset and long-term trends in the data. We emphasize that the cosmological analysis corresponds to data at higher frequencies ($>0.5$ Hz) and smaller spatial scales ($\ell\gtrsim 200$ or $\lesssim 1^{\circ}$) than those used here.
@@ 6.2 Mode separation and TOD selection
Figure 6 shows a typical receiver output for 40 minutes. The common mode signal in all three optics tubes and their bands is from the atmosphere and readily apparent. However, there are often optics tube-common modes that do not correspond to emission from the atmosphere or sky; these modes are periodic and depend on both scan position and direction, and thus clearly not on the sky. We attribute these to systematic pickup for each optics tube (hereafter “pickup”). When atmospheric fluctuations are weak, scan-synchronous pickup can become comparable to or even dominate atmospheric fluctuations, confounding analysis of the atmosphere. Binning and subtracting out a template of left- and right-going scans is common method of removing this pickup for CMB analysis, but it is less useful for atmospheric analysis because it may remove some of the atmosphere. We therefore want a method that separates atmospheric fluctuations and systematic pickup.
@@
where $g_{w,i}$ and $g_{p,i}$ are the responses of detector $i$ to the fluctuations in water vapor ($\delta w$) and pickup in optics tube $j$ ($\delta p_{j}$), respectively, and $\epsilon_{i}(t)$ is some residual. The pickup modes are common to an optics tube, with each detector in that tube having a different gain to it, [^10] Based on a phenomenological examination of the data, we find that a linear response of each detector to the underlying systematic mode in each optics tube is sufficient. while the atmosphere is common to all three optics tubes. [^11] Fluctuations smaller than ACT’s field of view are negligibly small compared to the common mode. With one common atmospheric mode and three optics tube pickup modes, we represent the ($n_{\text{detector}}\times n_{\text{time}}$) matrix of all detector fluctuations $\delta\mathbf{P}$ for each TOD, $t$, as
@@
where $\sigma^{2}_{\text{c}}$ and $\sigma^{2}_{\text{nc}}$ are the variances of the common and non-common modes in each detector’s signal. Distributions of $a^{2}$ for each band are shown in Figure B3. We set the threshold for a detector-TOD being “atmosphere-dominated” as $a^{2}\geq 10$, that is, when the variance of the atmospheric fluctuations is at least 10 times that of systematic fluctuations. We restrict the data to detector-TODs with this designation to minimize systematic contamination in quantifying the characteristics of atmospheric turbulence.
@@ 6.3 Structure functions
ACT employs a wide back-and-forth constant-elevation raster scan, meaning that the structure function describing its data has both a temporal and azimuthal dependence. To sample the temporal structure function, we take 256 logarithmically-spaced samples of $\tau$ between $0.1$ and $1000$ seconds and for each $\tau$ find all pairs of times $t_{i},t_{j}$ such that $t_{i}-t_{j}=\tau$ and $\phi(t_{i})=\phi(t_{j})$, where $\phi(t)$ is the azimuth of the boresight as a function of time. The structure function at $\tau$ is then the average of square differences $D(\tau)\langle[\delta P(t_{i})-\delta P(t_{j})]^{2}\rangle_{\textrm{pairs}}$. To probe the structure function at levels below the white noise of individual detectors, we take as $\delta P$ the average power of all detectors for each band.
@@ 6.4 Estimating the coefficient of turbulence and outer scale
We model the fluctuations, $y(t)$, as a zero-mean correlated multivariate Gaussian distribution. If the fluctuations are stationary, the covariance depends only on the delay $\tau=|t-t^{\prime}|$ and is given by Equation 6 as
@@ 7.1 Power spectrum, structure functions and outer scale estimates
Figure 7 shows the power spectra of the data in Figure 6. They agree with expectations after accounting for the effects of beam smearing. Evidence for the outer scale shows up at $f<0.01\,$Hz. Figure 8 shows the structure function for each band, averaged [^15] Because the scaling of the structure functions of different TODs can vary by several orders of magnitude, an explicit average will be most sensitive to TODs in which atmospheric fluctuations are the strongest. In order to probe the average shape of the structure function, we take a weighted average to normalize each structure function with respect to its variance. over all TODs that were designated as atmosphere-dominated ($a^{2}\geq 10)$. The key feature is the flattening for $\tau\gtrsim 100$s, beyond which there are no additional contributions to the variance. For each average structure function, we overlay a best fit of the structure function as predicted by the Matérn covariance and accounting for the beam profile and water vapor scale height (see below) as described in more detail in Appendix A.2. Figure 9 shows the distribution of $\tau_{0}$ from the fits for PA6 f150 which is representative of the other arrays. In the frozen transport approximation the spatial length scale differs from the temporal length scale by a factor of the wind speed. Thus, to find the outer scale $L_{0}$ we multiply by the wind speed derived from ERA5 (Hersbach et al., 2020). As shown in Morris et al. (2022), there is an excellent agreement between the wind speed derived from ACT with that derived from the absolute humidity weighted velocity profile in ERA5.
@@
We also show the distribution of the angular outer scale, computed by scaling $\tau_{0}$ by the angular velocity of the atmosphere estimated directly from ACT following the method in Morris et al. (2022). We see in Figure 9 that $L_{0}$ is comparable to the scale height of the water vapor. While not surprising, simultaneously accounting for both the scale height and outer scale leads to a somewhat modified form of the predicted structure function shown in Figure 8. We discuss this further in Appendix A.1.
@@ 7.2 Comparison of fluctuations in ACT to APEX
We find that both $\langle\delta P^{2}\rangle^{\frac{1}{2}}_{t}$ and $\langle\delta w_{z}^{2}\rangle^{\frac{1}{2}}_{t}$, where the subscript denotes average over a TOD, are distributed roughly log-normally, and hence the distribution of $g$ is log-normal. For a log-normal distribution, the maximum likelihood estimator of $g$ is the median (while the mean is upwardly biased for any nonzero variance due to a heavy tail). To calibrate ACT’s fluctuation to APEX, we account for the difference in height between ACT and APEX ($\sigma_{ACT}=0.94\,\sigma_{APEX}$, Section 4). This results in a scaled value for $g$ that we call $g_{s}$. To estimate ACT’s efficiency, $\eta$, we divide $g_{s}$ by the entry for $\partial P/\partial w_{z}$ in Table C1. Because the resolution of APEX is 0.01 mm, we compute the correlation for RMS fluctuation levels of 0.02 mm and above. As this cutoff level is increased, the computed $\eta$ decreases. The results are reported in Table 1. Based on the variance over seasons, the dispersion of the correlation in Figure 11, the bias associated with the cutoff level, and the uncertainty in the passbands, we estimate the net uncertainty to be roughly 20%.
@@
There is an overall degeneracy, that we cannot determine, between $\eta$ and any out-of-band passband leakage or an incorrect passband measurement. The quoted values are for the recommended frequency bounds for the integrals over the passbands (Hasselfield et al., 2024) and is based on an analysis of the Fourier Transform Spectrometer (FTS) used to measure them Alford (2024). For the 90 GHz bands, there is possible band pass leakage at the 1% level in the FTS measurements that resembles a second harmonic of the primary passband. Were it real as opposed to a measurement artifact, it would lead to sensitivity to the large atmospheric signal near 180 GHz shown in Figure 4. However, there is no clear evidence of sensitivity to this feature.